OpenStudy (anonymous):

mr.cruz made a swimming pool infront of their house measuring 10m by 6m by 5m. how much did he get from the sale of the soil dug up if a cubic meter of soil was sold at 50 dollars

OpenStudy (nubeer):

vol = 10*6*5 vol = 300m^3 1m^3 = 50$ 300m^3 = 50*300 Price = 15000$

OpenStudy (anonymous):

thank you

OpenStudy (nubeer):

you're welcome